OpenStudy (anonymous):

Help please if slope 3 , y-interccept -2 what is a equation of the line

OpenStudy (trojanpoem):

y = mx + b ( b: y-intercept, m : slope) Can you continue ?

OpenStudy (anonymous):

i found the answer already thanks though i do have another one though

OpenStudy (anonymous):

Find the slope of the line containing the pair of points (8,2) and (9,9)

OpenStudy (trojanpoem):

y2 - y1 / x2 - x1 = slope

OpenStudy (anonymous):

ok so 9-2 = 7 9-8 = 1 7/1 is my answer

OpenStudy (anonymous):

Right

OpenStudy (trojanpoem):

9 - 2/ 9 -8 = 7/1 = 7 You' re correct !